This incident has been resolved.
Jun 23, 23:12 UTC
All systems are stabilized and we are continuing to monitor for the next few hours.
Jun 23, 21:31 UTC
BGP telemetry has been fully restored and we are continuing to monitor all systems and services.
Jun 23, 19:10 UTC
v6 BGP telemetry has been restored
Jun 23, 18:45 UTC
During stabilization, we had to restart some services, resulting in another 10 minute drop in connectivity from approximately 1750UTC to 1800 UTC. Connectivity should be stable now and we are continuing to monitor. BGP telemetry remains down and we are working to restore.
Jun 23, 18:02 UTC
General connectivity for portal, APIs, and ingest has been restored and we are continuing to monitor and work to stabilize the platform. BGP sessions remain down.
Jun 23, 17:16 UTC
The issue has been identified and we are working to resolve it.
Jun 23, 17:10 UTC
Access to web portal, APIs, ingest, and bgp peering have been affected.
Jun 23, 17:06 UTC